A movie rental co has 2 plans. One charges $7 per month and $1 per movie rented. The other one charges $20 per moth and $.50 per movie. Bill has the first plan and Deanna has the second plan. Last month they both paid the exact same amount and rented the exact same number of movies. Exactly how much did they each pay ?
Answer by richwmiller(17219)

y=mx+b
y=x+7
y=x/2+20
x+7=x/2+20
x=x/2+13
x/2=13
x=26
They rented 26 movies.
They each paid 33 dollars